# Latent Multimodal Reconstruction for Misinformation Detection

### v1

@@ 6.1 Training Datasets
For the “True vs. MC” experiments, we use the four versions of our “MisCaption This!” dataset, as detailed in Section 4, along with two additional datasets: the deduplicated version of the Crossmodal HArd Synthetic MisAlignment (CHASMA) dataset, which includes 145,891 truthful and 145,891 miscaptioned images [28], and the CLIP-based Named Entity Swapping by Topic (NESt) dataset, containing 847,693 miscaptioned images and 1,007,744 truthful images [17]. For “True vs. OOC” experiments, we use the NewsCLIPpings dataset [14], Merged/Balanced version, comprising 42,680 truthful and OOC samples in total. Finally, for multi-class classification, we combine one of the “miscaptioned images” datasets (“MisCaption This!”, CHASMA, or NESt) with the NewsCLIPpings dataset, which represents the OOC class, and apply random under-sampling to balance the classes.
@@ 6.2 Evaluation Protocol
We adhere to the training, validation, and testing splits provided by each dataset and evaluate models using Accuracy as the primary metric. After training on any of the aforementioned synthetic datasets, we further assess performance on the VERITE benchmark [28], which comprises 1,000 real-world samples: 338 truthful image-caption pairs, 338 miscaptioned images, and 324 out-of-context pairs. For binary classification, we report “True vs. OOC” and “True vs. MC” accuracies, while for the multi-class task, we provide overall accuracy along with per-class accuracy. All three tasks serve as out-of-distribution evaluations, as the training data are synthetically generated, while the final evaluation is conducted on real-world data.
@@ 6.3 Competing Methods
In addition to our proposed LAMAR architecture and its variations, we reproduce and evaluate the performance of several competing methods, without using any external evidence: (1) DT-Transformer [17]: a Transformer encoder that processes the input sequence $[\mathbf{CLS};\mathbf{I};\mathbf{C}]$; (2) RED-DOT (‘Baseline’ version) [21]: a Transformer-based model utilizing the fused representation $[\mathbf{CLS};\mathbf{F}]$ as input; (3) AITR (Attentive Intermediate Transformer Representations) [43], a model incorporating multimodal similarity (MUSE) and a self-attention mechanism over a stack of Transformer encoder blocks, with varying numbers of multi-head self-attention layers. We consider both its “attention pooling” and “weighted pooling” variants.
@@ 6.4 Implementation Details
We train LAMAR using a Transformer encoder with 4 encoder blocks, each containing 4 multi-head self-attention heads and a feed-forward layer of 1,024 dimensions and a dropout of 0.1 probability. The model is optimized with Adam, using a learning rate of $41e-$ 4 and a batch size of 512. We train the network for up to 50 epochs, with early stopping after 10 epochs if validation performance does not improve. For the large-scale pre-training, we consider $\sigma\in\{0.1,0.2\}$ and $\mu=0.0$ and dropout probability $dp\in\{0.2,0.5\}$. To ensure reproducibility, we set a constant random seed of 0 for PyTorch, Python Random, and NumPy.
@@ 6.5 Computational Complexity
Generating a single version of “MisCaption This!”  including the generation of captions for 35,536 images, took 18.3 hours using a single Nvidia GeForce RTX 4090 (24GB RAM). Feature extraction with CLIP ViT-L/14 required approximately 16 minutes for images and 2 minutes for texts, using a batch size of 256. Both data generation and feature extraction were repeated four times for each version of “MisCaption This!”.
Using fvcore [^3] https://github.com/facebookresearch/fvcore, we estimate that LAMAR, trained end-to-end with a gate mechanism, requires 980,997 FLOPs, and 1,034,181 FLOPs with the attention mechanism. This is comparable to AITR (1,070,653 FLOPs) [43] and significantly more efficient than RED-DOT (2,208,072 FLOPs) [21]. When using pre-extracted features, LAMAR requires a maximum of 30 seconds per epoch (batch size 512) to process the full “MisCaption This!” for multiclass classification.
Figure 4: Performance of detection models (DT-Transformer and RED-DOT) trained on four variations of the “MisCaption This!” dataset ($\mathcal{D}_{1}$, $\mathcal{D}_{2}$, $\mathcal{D}_{3}$, and $\mathcal{D}_{4}$), evaluated under varying filtering thresholds ($l\in{0,5,10,15,25,50,\text{None}}$) in terms of test-set accuracy and VERITE “True vs. MC” accuracy.
@@ 7.1 Dataset Variants and Filtering
We first examine the generalizability of two established MMD models: DT-Transformer [17] and RED-DOT [21] (without incorporating external evidence) when trained on the four “MisCaption This!” variations ($\mathcal{D}_{1}$, $\mathcal{D}_{2}$, $\mathcal{D}_{3}$, and $\mathcal{D}_{4}$), under different filtering thresholds $l\in\{0,5,10,15,25,50,\text{None}\}$. As shown in Fig. 4, test-set accuracy tends to improve with higher $l$ values across all datasets and models. The highest performance is observed with no filtering ($l=\text{None}$), ranging from 79% to 83.4%. In contrast, evaluating models on the real-world data of VERITE (“True vs. MC”), there is a slight negative (Pearson) correlation between test-set accuracy and VERITE accuracy: -0.29 for DT-Transformer and -0.17 for RED-DOT. This suggests that higher test-set accuracy does not guarantee better generalization, as training with synthetic misinformation that is too easily distinguishable from truthful pairs may lead to inflated test-set accuracy without improving performance on real-world misinformation.
Both models achieve peak performance at $l=10$, with accuracies of 63.3% and 61.8%, followed by $l=5,15,$ and $25$; highlighting the importance of filtering out “rambling” and generic descriptions sometimes produced by LLaVa. Thus, by refining the dataset in this way, we improve model robustness and enhance real-world generalization on VERITE. Overall, both models achieve their highest performance when trained on $\mathcal{D}_{3}$, with RED-DOT and DT-Transformer averaging 60.2% and 59.2%, respectively. This suggests that $p_{3}^{gen}$ generates misinformation patterns that more closely resemble real-world examples, leading to better model generalization. Therefore, for subsequent experiments, we will continue using the $\mathcal{D}_{3}$ version of “MisCaption This!”, while treating $l$ as a tunable hyper-parameter.
Table 1: Comparative and ablation analysis of the proposed LAMAR method and prior SotA models (DT-Transformer, RED-DOT, and AITR) trained on three synthetic datasets (“MisCaption This!” ($\mathcal{D}_{3}$), NESt, and CHASMA) and evaluated on VERITE (“True vs. MC”). Bold denotes the highest accuracy on VERITE.
Model $\mathcal{D}_{3}$ NESt CHASMA
DT-Transformer 61.8 58.7 57.8
RED-DOT 63.3 57.0 58.0
AITR 62.9 58.4 58.3
LAMAR Variant $\mathcal{D}_{3}$ NESt CHASMA
@@
DT-Transformer 79.7 69.4
RED-DOT 81.5 73.5
AITR 84.1 74.1
@@
7.2 Ablation across Datasets
We train three established MMD models, DT-Transformer, RED-DOT, and AITR, alongside various variants and ablations of the proposed LAMAR method using three synthetic “True vs. MC” training datasets: “MisCaption This!” ($\mathcal{D}_{3}$), CHASMA [28], and NESt [17], and evaluating their performance on VERITE (“True vs. MC”).
As shown in Table 1, using “MisCaption This!” ($\mathcal{D}_{3}$) as the training dataset consistently results in models with higher generalizability to real-world data (VERITE). This trend holds across all evaluated models and variants. These results support our hypothesis that LVLMs are capable of generating more effective synthetic training data, ultimately leading to more robust detection models, compared to previous approaches that rely on named entity swaps or cross-modal misalignment. Furthermore, we observe that LAMAR trained end-to-end (E2E) with the “Gate” mechanism for integration achieves the highest overall performance on VERITE across all three datasets, yielding the best performance (66%) when using the “MisCaption This!” ($\mathcal{D}_{3}$); achieving 7.8% higher performance compared to when trained on NESt, and 10.4% higher compared to CHASMA. Overall, we observe that E2E training tends to yield better performance, as it is directly optimized to reconstruct the embeddings of actual miscaptioned images, in contrast to PT methods that reconstruct embeddings of truthful pairs interjected with Gaussian noise or dropout.
Additionally, LAMAR with E2E training, utilizing either the “Gate” or “Attention” mechanisms for integrating the reconstruction embeddings, consistently outperforms the direct integration; emphasizing the importance of utilizing an integration mechanism. The “Mask” mechanism does not perform as well, likely due to the complete masking of the reconstruction embedding, which reduces its expressiveness. Instead, modulating the embedding by adjusting its values based on learned gate values or attention scores proves to be a more effective strategies.
We also investigate unimodal reconstruction as an ablation. Direct integration without the image input (Direct-Image), using only the text as input, results in a slight decrease in performance (-0.9%). Similarly, excluding textual information (Direct-Text), which effectively transforms the task into a latent image captioning approach, leads to even lower performance (-1.3%). Removing the reconstruction network entirely from LAMAR, leaving only the detection classifier, yields the lowest performance among all end-to-end (E2E) approaches. Finally, we observe that LAMAR [E2E, Gate] outperforms prior state-of-the-art models, such as DT-Transformer, RED-DOT, and AITR, by 6.8%, 4.3%, and 4.9%, respectively, when trained on “MisCaption This!” ($\mathcal{D}_{3}$); validating the effectiveness of our proposed method for the detection of miscaptioned images.
@@ 7.2 Ablation across Datasets
Table 3: Performance of models trained on the multi-class “MisCaption This!” ($\mathcal{D}_{3}$) or the combined CHASMA and NewsCLIPpings datasets, evaluated on VERITE, reported as overall accuracy and per-class accuracy.
Training Model Accuracy True MC OOC
CHASMA DT-Transformer 50.0 78.7 23.1 48.0
RED-DOT 48.5 71.9 21.9 51.7
AITR 51.4 90.5 15.4 48.0
LAMAR [E2E, Gate] 53.2 89.6 18.1 51.7
$\mathcal{D}_{3}$ DT-Transformer 47.8 53.0 60.4 29.2
RED-DOT 48.8 50.9 53.3 41.9
AITR 51.7 50.3 62.7 41.5
LAMAR [E2E, Gate] 54.2 58.6 58.6 44.9
7.3 Out-of-context and Multiclass Detection
@@
Table 2 presents the performance of LAMAR compared to prior SotA models on the binary classification task of “True vs. OOC”, without external evidence. LAMAR achieves the highest accuracy on the NewsCLIPpings dataset, with LAMAR [E2E, Attention] reaching 84.8% and LAMAR [E2E, Gate] closely following at 84.7%, outperforming AITR (84.1%) and other models. More importantly, when evaluated on VERITE, LAMAR [E2E, Gate] attains 76.3%, surpassing the best prior model (AITR) by 3%. We observe a noticeable performance gap between “True vs. OOC" (76.3%) and “True vs. MC" (66.0%) due to differing task complexity. In OOC, the entire image is mismatched with the caption, making discrepancies more apparent, while MC involves subtle manipulations (e.g., actions, dates, locations, names), making it harder to distinguish from truthful captions.
For the multi-class classification task, we consider two training datasets: “MisCaption This!” ($\mathcal{D}_{3}$) and CHASMA combined with NewsCLIPpings. As shown in Table 3, LAMAR [E2E, Gate] continues to demonstrate superior generalization, regardless of the dataset used for training. When trained on “MisCaption This!” ($\mathcal{D}_{3}$), LAMAR achieves the highest overall accuracy (54.2%), representing a relative improvement of 4.8%, 11.1%, and 13.4% over AITR, RED-DOT, and DT-Transformer, respectively. Notably, all four models achieve relatively higher accuracy on the MC class when trained on “MisCaption This!” ($\mathcal{D}_{3}$) compared to CHASMA+NewsCLIPpings. These results validate the effectiveness of our proposed method in leveraging synthetic data generated by an LVLM and highlight the impact of incorporating a reconstruction network to enhance the detection process, ultimately improving generalization to real-world misinformation detection.
8 Conclusions

### v3

@@ 6.1 Training Datasets
For the “True vs. MC” experiments, we use the 4 versions produced with LLaVA, 1 from Llama, 1 from Molmo, as detailed in Section 4, along with two additional datasets: the deduplicated version of the Crossmodal HArd Synthetic MisAlignment (CHASMA) dataset, which includes 145,891 truthful and 145,891 miscaptioned images [28], and the CLIP-based Named Entity Swapping by Topic (NESt) dataset, containing 847,693 miscaptioned images and 1,007,744 truthful images [30]. For “True vs. OOC” experiments, we use the NewsCLIPpings dataset [22], Merged/Balanced version, comprising 42,680 truthful and OOC samples in total. Finally, for multi-class classification, we combine one of the “miscaptioned images” datasets (“MisCaption This!”, CHASMA, or NESt) with the NewsCLIPpings dataset, which represents the OOC class, and apply random under-sampling to balance the classes.
@@ 6.2 Evaluation Protocol
We adhere to the training, validation, and testing splits provided by each dataset and evaluate models using Accuracy as the primary metric. After training on any of the aforementioned synthetic datasets, we further assess performance on the VERITE benchmark [28], which comprises 1,000 real-world samples: 338 truthful image-caption pairs, 338 miscaptioned images, and 324 out-of-context pairs. We report ‘True vs. OOC’ and ‘True vs. MC’ accuracies for binary classification tasks, and overall accuracy for the multi-class task. All three tasks serve as out-of-distribution evaluations, as the training data are synthetically generated, while the final evaluation is conducted on real-world data.
For the ‘True vs. MC’ task, we update the VERITE benchmark, following the data selection criteria and labeling process defined in [28]. Specifically, we collected 50 new samples – sourced from Snopes and Reuters, balanced across the two classes – describing events that occurred between October 2024 and July 2025. Crucially, these data points fall after the knowledge cut-off point of the selected VLMs; Molmo, Llama, and LLaVA. This expansion, which we denote VERITE 24/25, was collected to examine the out-of-distribution robustness and temporal generalization capacity of both zero-shot VLMs and our proposed methods against emerging misinformation.
@@ 6.3 Competing Methods
In addition to our proposed LAMAR architecture and its variations, we reproduce and evaluate the performance of several competing methods: (1) MUSE [29]: a simple similarity-based baseline consisting of a single hidden layer applied to cosine similarity scores between image-text pairs. (2) SpotFake* [34]: A neural network that processes the input sequence $\mathbf{I};\mathbf{C}]$. (3) SAFE* (Similarity-Aware FakE) [43]: Similar to SpotFake*, but processing the input sequence $\mathbf{I};\mathbf{C};sim(\mathbf{I},\mathbf{C})]$. (4) MVAE* (Multimodal Variational Autoencoder) [17]: A bimodal variational autoencoder that maps image and text features into a shared latent space. The network is trained to jointly minimize the reconstruction, KL divergence, and cross-entropy losses. The decoder reconstructs the original input words and the image embeddings from the shared latent space. (5) BCMF* (Bidirectional Cross-Modal Fusion) [40]: A dual-stream architecture that performs bidirectional cross-attention between images and captions, followed by gating mechanisms to refine the features. The final fused feature vector concatenates the refined features and their explicit interaction terms, expressed as $\text{Maxpool}(\mathbf{C}^{\prime});\mathbf{I}^{\prime};|\text{Maxpool}(\mathbf{C}^{\prime})-\mathbf{I}^{\prime}|]$, which is then used for the final classification. (6) PMFN* (Progressive MLP-Mixer Fusion Network) [16]: A multi-stage architecture, that performs hierarchical fusion. It leverages the MLP-Mixer module to iteratively refine multimodal features by fusing image and text representations from different abstraction stages of the backbones, aiming to enhance the correlation between shallow and deep features before final classification. (7) DT-Transformer [30]: a Transformer encoder that processes the input sequence $\mathbf{CLS};\mathbf{I};\mathbf{C}]$. (8) RED-DOT [27]: a Transformer-based model operating on the fused representation $\mathbf{CLS};\mathbf{F}]$; we use the Baseline version without external evidence. (9) AITR (Attentive Intermediate Transformer Representations) [29]: A model that uses a stack of $n$ Transformer encoder layers of varying number of attention heads to process the initial multimodal sequence $\mathbf{CLS}_{0};\mathbf{F}]$ and then applies a self-attention mechanism over the sequence of intermediate classification tokens $\mathbf{CLS}_{1},\dots,\mathbf{CLS}_{n}]$ for final classification.
For a fair comparison, all models marked with an asterisk (*) were adapted to use CLIP ViT-L/14 image and text features, replacing their original backbones (VGG-19, BERT, DeiT, Swin Transformer, Bi-LSTM, etc.).
@@ 6.4 Implementation Details
We train LAMAR using a Transformer encoder with 4 encoder blocks, each containing 4 multi-head self-attention heads and a feed-forward layer of 1,024 dimensions and a dropout of 0.1 probability. The model is optimized with Adam, using a learning rate of 1e41e-4 and a batch size of 512. We train the network for up to 50 epochs, with early stopping after 10 epochs if validation performance does not improve. For the large-scale pre-training, we consider $\sigma\in\{0.1,0.2\}$ and $\mu=0.0$ and dropout probability $dp\in\{0.2,0.5\}$. To ensure reproducibility, we set a constant random seed (0) for PyTorch, Python Random, and NumPy.
@@ 6.5 Computational Complexity
Generating a “MisCaption This!” dataset with LLaVA, took 18.3 hours using a single Nvidia GeForce RTX 4090 (24GB RAM). Feature extraction with CLIP ViT-L/14 required approximately 16 minutes for images and 2 minutes for texts, using a batch size of 256; and was performance only once per dataset, and is used for any model.
Using fvcore [^5] https://github.com/facebookresearch/fvcore, we estimate that LAMAR, trained end-to-end with a gate mechanism, requires 980,997 FLOPs, and 1,034,181 FLOPs with the attention mechanism. This is comparable to AITR (1,070,653 FLOPs) [29] and more efficient than RED-DOT (2,208,072 FLOPs) [27]. When using pre-extracted features, LAMAR requires a maximum of 30 seconds per epoch (batch size 512) to process the unfiltered ($l=None$) dataset for multiclass classification.
@@ 7 Results
Figure 5: Performance of detection models (DT-Transformer and RED-DOT) trained on four variations of LLaVA-$\mathcal{D}_{1}$, $\mathcal{D}_{2}$, $\mathcal{D}_{3}$, and $\mathcal{D}_{4}$, evaluated under varying filtering thresholds ($l\in\{0,5,10,15,25,50,\text{None}\}$, or 4.5%, 19.1%, 27.8%, 34.9%, 47.9%, 74.0%, and 100% of the dataset) in terms of Test-set Accuracy and VERITE “True vs. MC” Accuracy.
@@ 7.1 Dataset Variants and Filtering
First, we investigate how different synthetic data variations ($\mathcal{D}_{1}$$\mathcal{D}_{4}$) and filtering thresholds ($l$) influence a model’s ability to detect synthetic and real-world misinformation. Specifically, we analyze the relationship between internal (synthetic) test-set performance and out-of-distribution generalization to the VERITE (“True vs. MC”) benchmark. As shown in Fig. 5, test-set accuracy tends to improve with higher $l$ values across all datasets and two established MMD models, DT-Transformer [30] and RED-DOT [27]. The highest test-set performance is observed without any filtering ($l=\text{None}$), ranging from 79% to 83.4%. However, we identify a divergence between test-set and real-world performance: with a negative Pearson correlation of 0.29-0.29 for DT-Transformer and 0.17-0.17 for RED-DOT. This indicates that high test-set accuracy on synthetic data does not guarantee better generalization; rather, training on synthetic misinformation that is too easily distinguishable from truthful pairs likely encourages the model to exploit spurious shortcuts rather than learning robust factual consistency, ultimately hindering its generalization.
Regarding filtering, both models achieve peak performance at $l=10$, with accuracies of 63.3% and 61.8%, followed by $l=5$, $15$, and $25$. This highlights the importance of filtering out the “rambling” or generic descriptions occasionally produced by VLMs. Moreover, filtering overly long texts prevents the detection model from relying on superficial length-based patterns (e.g., that longer texts tend to be “falsified”) rather than actual factual consistency. Thus, by refining the dataset in this way, we improve model robustness and enhance real-world generalization on VERITE.
Overall, both models achieve their highest performance when trained on $\mathcal{D}_{3}$, with RED-DOT and DT-Transformer averaging 60.2% and 59.2%, respectively. This suggests that $p_{3}^{gen}$ produces misinformation patterns that more closely resemble real-world examples, resulting in better model generalization. Consequently, we adopt $p_{3}^{gen}$ to produce Molmo-$\mathcal{D}_{3}$ and Llama-$\mathcal{D}_{3}$, and use LLaVA-$\mathcal{D}_{3}$ variant of “MisCaption This!” for proceeding experiments, while treating $l$ as a hyperparameter.
7.2 Ablation across Datasets
Table 1: LAMAR ablations trained on three synthetic datasets (LLaVA-$\mathcal{D}_{3}$, NESt, or CHASMA) and evaluated on VERITE (“True vs. MC”). Bold denotes the highest accuracy.
LAMAR Variant LLaVA-$\mathcal{D}_{3}$ NESt CHASMA
@@ 7.2 Ablation across Datasets
No Reconstruction 62.7 58.2 56.7
We conduct an extensive analysis of the proposed LAMAR method, including various model variants and ablations. This analysis is conducted using three distinct synthetic datasets for training: “MisCaption This!” (LLaVA-$\mathcal{D}_{3}$), CHASMA [28], and NESt [30]. The performance of all models is consistently evaluated on the VERITE benchmark, specifically, the “True vs. MC” task.
As shown in Table 1, training on “MisCaption This!” (LLaVA-$\mathcal{D}_{3}$) consistently yields models with stronger generalization to real-world data. This trend holds across all LAMAR variants and supports our hypothesis that VLM-generated training data provides more effective supervision than earlier methods based on named entity swaps or cross-modal inconsistencies. The best overall performance (66%) is achieved by LAMAR [E2E] with the Gate integration mechanism trained on “MisCaption This!” data (LLaVA-$\mathcal{D}_{3}$), outperforming NESt by 7.8% and CHASMA by 10.4%. E2E training proves advantageous, as it directly optimizes reconstruction over actual miscaptioned pairs, unlike PT variants that rely on perturbed truthful captions via noise or dropout. These results highlight the value of semantically grounded, VLM-generated training data that more closely resembles real-world misinformation.
Among integration strategies, Gate and Attention mechanisms consistently outperform direct embedding injection, highlighting the importance of adaptive integration. In contrast, the Mask mechanism underperforms, likely due to reduced representational capacity from fully masking the reconstruction embedding. Similarly, removing the fusion component $\mathbf{F}$ (Eq. 4) and using simple concatenation ($\mathbf{I};\mathbf{C}]$) significantly degrades performance, emphasizing the role of effective multimodal fusion.
We also conduct unimodal ablations. Removing the image input (Direct-Image) leads to a slight drop in performance (-0.9%), while removing the text (Direct-Text) results in a larger drop (-1.3%), effectively turning the task into latent captioning. Finally, eliminating the reconstruction module entirely yields the lowest E2E performance, confirming the critical role of the reconstruction in guiding accurate detection.
To evaluate how LAMAR integrates the reconstructed signal with the primary inputs, we analyze the distribution of attention scores across input tokens (image, caption, and reconstructed caption embeddings) and the behavior of the gating mechanism. Attention scores are substantially higher on the reconstruction token compared to the image or input caption tokens across classes. Truthful pairs receive higher attention on average for both the image token (0.218 vs. 0.176) and the input caption (0.213 vs. 0.169), while falsified pairs shift attention toward the reconstructed caption (0.655 vs. 0.569). Similarly, the gate score correlates positively with reconstruction loss (0.451), and reconstruction loss itself is positively associated with the model’s predicted label (0.386), suggesting that higher reconstruction error pushes the model toward a “falsified” prediction. These patterns indicate that LAMAR learns to focus more on the reconstructed caption, via attention or gating, when detecting inconsistencies, making it a valuable auxiliary signal for accurate classification. Thus, when the reconstruction caption embeddings diverges semantically from the input embeddings, the model flags the pair as falsified.
7.3 Comparative Study across VLM-Generated Datasets
Table 2 shows the performance of LAMAR and prior SotA models trained on VLM-generated datasets produced by LLaVA, Molmo, and Llama, and evaluated on the VERITE benchmark (“True vs. MC”). Across all three datasets, LAMAR consistently outperforms prior models by significant margins: 14.4% over MUSE, 8.7% over MVAE, 7.0% over SpotFake, 6.3% over SAFE, 3.8% over BCMF, 3.8% over MPFN, 6.8% over DT-Transformer, 4.3% over RED-DOT, and 4.9% over AITR. This holds for both LAMAR variants using gated or attention-based integration. The best overall result is achieved by LAMAR [E2E, Gate] trained on LLaVA-$\mathcal{D}_{3}$, reaching 66.0% accuracy.
Table 2: Performance of models trained on three VLM-generated datasets (LLaVA-$\mathcal{D}_{3}$, Molmo-$\mathcal{D}_{3}$, Llama-$\mathcal{D}_{3}$) and evaluated on VERITE (“True vs. MC”). Bold indicates the highest overall accuracy; Underline indicates the best accuracy per training dataset.
Model LLaVA-$\mathcal{D}_{3}$ Molmo-$\mathcal{D}_{3}$ Llama-$\mathcal{D}_{3}$
MUSE [29] 57.7 54.9 57.5
MVAE [17] 60.7 60.8 60.5
SpotFake [34] 61.7 62.6 60.7
SAFE [43] 62.1 63.0 59.6
BCMF [40] 63.6 61.2 59.0
MPFN [16] 63.6 64.1 58.6
DT-Transformer [30] 61.8 62.1 58.4
RED-DOT [27] 63.3 62.9 61.7
AITR [29] 62.9 63.2 63.0
LAMAR [E2E, Attention] 65.1 65.1 64.2
LAMAR [E2E, Gate] 66.0 64.4 63.9
LLaVA-generated data yielding higher performance may be partially attributed to the use of “adversarial prompt selection” on it, whereas Molmo and Llama data were generated using the same prompt ($p_{3}^{gen}$). While this introduces a slight advantage for LLaVA, it is notable that models trained on Molmo- and Llama-generated data performed comparably well, thus underscoring the general utility of VLM-generated data across models.
Qualitative analysis (see Fig. 3) further highlights that LLaVA-$\mathcal{D}_{3}$ contain more nuanced and believable false captions by maintaining semantic consistency with the image. In contrast, LLaVA-$\mathcal{D}_{4}$ and Llama-$\mathcal{D}_{3}$ often contain more apparently false captions – for example, attributing protests to “Madrid, Spain” or “in front of the Eiffel Tower in Paris” despite clear visual cues such as Greek flags. Such easily detectable falsehoods are less reflective of real-world misinformation and thus less effective for training models that generalize well.
7.4 Out-of-context and Multiclass Detection
@@
MVAE [17] 60.6 55.2
MUSE [29] 80.6 72.0
SpotFake [22] 79.8 67.1
SAFE [43] 83.7 72.7
BCMF [40] 76.6 67.1
MPFN [16] 76.1 66.5
DT-Transformer [30] 79.7 69.4
RED-DOT [27] 81.5 73.5
AITR [29] 84.1 74.1
@@
Table 4: Comparison of multi-class models trained either on CHASMA, LLaVA-$\mathcal{D}_{3}$, Molmo-$\mathcal{D}_{3}$, and evaluated on VERITE.
Model CHASMA LLaVA-$\mathcal{D}_{3}$ Molmo-$\mathcal{D}_{3}$
MUSE [29] 38.5 49.7 47.3
MVAE [17] 39.8 40.5 42.7
SpotFake [34] 49.2 48.7 47.2
SAFE [43] 50.4 51.4 51.0
BCMF [40] 45.9 45.4 47.5
MPFN [16] 46.1 44.1 44.8
DT-Transformer [30] 50.0 47.8 48.7
RED-DOT [27] 48.5 48.8 49.1
AITR [29] 51.4 51.7 52.0
LAMAR [E2E, Gate] 53.2 54.2 54.9
@@
Table 3 presents LAMAR’s performance compared to prior SotA models on “True vs. OOC” classification, without access to external evidence. LAMAR achieves the highest accuracy on the NewsCLIPpings dataset, with [E2E, Attention] reaching 84.8% and [E2E, Gate] closely following at 84.7%, both outperforming AITR (84.1%) and other models. More importantly, LAMAR [E2E, Gate] achieves 76.3% on VERITE, surpassing the best prior model (AITR) by 3%. We observe a consistent performance gap between True vs. OOC” (76.3%) and True vs. MC” (66.0%), reflecting the greater difficulty of detecting subtle manipulations (e.g., dates, places, or actions) in MC cases, compared to the more blatant mismatches typical of OOC examples.
For the multi-class classification task, we employ three training datasets: LLaVA-$\mathcal{D}_{3}$, Molmo-$\mathcal{D}_{3}$, and CHASMA; all being combined with NewsCLIPpings to represent the OOC class. As shown in Table 4, LAMAR [E2E, Gate] consistently generalizes best across all training sets. When trained on Molmo-$\mathcal{D}_{3}$, LAMAR achieves the highest overall accuracy (54.9%), representing a relative improvement of 5.6% over AITR. These results further validate the benefits of using VLM-generated data in LAMAR.
Table 5: Comparison of zero-shot VLMs against fine-tuned LAMAR variants on the “True vs. MC” task of VERITE and the post-knowledge-cutoff data of VERITE 24/25.
Model Parameters VERITE VERITE 24/25
Llama 3.2 11B 59.3 53.1
Molmo 7B 63.0 58.0
LLaVA 1.6 7B 66.9 62.0
LAMAR [E2E, Attention] 21M 65.1 70.0
LAMAR [E2E, Gate] 19M 66.0 68.0
7.5 Temporal Generalization Analysis
Table 5 compares the performance of the three VLMs used for the generation of “MisCaption This!” data, in a zero-shot setting against our fine-tuned LAMAR variants on the VERITE “True vs. MC” task. On the original VERITE dataset, the performance of the VLMs varies, with LLaVA 1.6 (66.9%) showing comparable performance to LAMAR [E2E Gate] (66.0%) and outperforming Molmo (63.0%) and Llama 3.2 (59.3%). Decent zero-shot performance is not unexpected, as these VLMs are trained on vast, large-scale datasets, making it highly probable that they have encountered samples or related content from popular fact-checking sources like Snopes and Reuters used in the original VERITE benchmark.
However, the analysis shifts when evaluated on the VERITE 24/25 set comprising samples dating after the VLMs’ knowledge cut-off. Here, all zero-shot VLMs demonstrate a notable decline in performance, with LLaVA 1.6 dropping to 62.0%, Llama 3.2 falling to 53.1%, and Molmo to 58.0%. In contrast, both fine-tuned LAMAR variants demonstrate strong temporal generalization, with LAMAR [E2E Attention] achieving 70.0% and LAMAR [E2E Gated] 68.0% on the post-cutoff data, despite having only $\approx$20 million parameters compared to the 7-11 billion parameters of the VLMs. The performance gap on VERITE 24/25 highlights that VLM’s zero-shot performance is limited by their training data and knowledge cutoff, whereas LAMAR’s reconstruction-based approach provided higher robustness against emerging misinformation.
8 Ethical Discussion
Our work demonstrates the potential of VLMs for generating high quality training datasets for multimodal misinformation detection. Nevertheless, it is crucial to consider the ethical implications of this work and future research in this direction. In our initial attempts, earlier open-source VLMs such as MiniGPT and LLaVA 1.5 struggled to generate coherent or realistic false captions. However, later versions (LLaVA 1.6, Llama 3.2, and Molmo) were more capable of producing false but plausible captions. While this capability proved valuable for research – enabling the creation of more robust training datasets – it also raises serious societal concerns as even lightweight, open-source models can now be repurposed to automate and scale disinformation campaigns. We believe that the research community must remain vigilant, ensuring that advancements in detection do not inadvertently provide a blueprint for malicious generation. To mitigate these risks, we refrain from publicly releasing specific prompts used to generate the “MisCaption This!” data. However, to ensure scientific reproducibility, we provide the general templates within this work, and the exact prompts will be made available to verified researchers upon reasonable request for non-commercial research purposes.
9 Conclusion